Session: Feb 2006 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click

Experiment Feedback

Medicina (Morsiani at Sat Mar 4 11:01:47 2006) Minor failures :

Urumqi (Liu Xiang at Sat Mar 4 11:11:51 2006) Minor failures : FORM
Formatter problem from the start to day062 UT13:54, from day062 UT16:05 to 16:33 the 5MHz (to formatter) problem. The problems were resolved, the recording are ok for left time.

Shanghai (TAO AN at Sat Mar 4 11:58:35 2006) Minor failures : RECO
we changed the bank from B to A at UT 12:12:00 because of the disk problem ,so we lost the first 10 scans.

Hartebeesthoek (Jonathan Quick at Sun Mar 5 11:39:58 2006) Minor failures : WIND RFI
Persitant spike at 132.5 MHz in IF especially RCP, otherwise repeated bursts of broadband noise. Wind stowed from 062.22:19-22:43, 22:49-22:54, 23:06-23:11 and 063.05:25-05:36 UT

Onsala (Michael Lindqvist at Sun Mar 5 12:50:16 2006) Success :
RFI seen in the bandpass. Otherwise no known problems.

Torun (Kaz Borkowski at Mon Mar 6 06:31:54 2006) Minor failures :
6-min. scan starting at 00:14 not recorded during disk-pack change. Otherwise no known problems.

Westerbork (Tony Foley at Mon Mar 6 13:04:31 2006) Minor failures :
Wb used in phased array mode. A little snow, otherwise OK

Jodrell_Bank (Alastair Gunn at Mon Mar 6 13:07:09 2006) Minor failures : ANT
The Lovell Telescope was used (Jb1). A local network problem at JBO led to the telescope control computer resetting its clock frequently during this experiment. Problem started at about 11:06 UT and continued for most of that day, after which problem was intermittent. This resulted in frequent periods of incorrect source tracking, lasting up to 0.5 hours, at intervals of up to 1 hour. Expect fringe loss during these periods.

Noto (P. Cassaro at Tue Mar 7 16:13:33 2006) Minor failures : RECO
No recording from 02:15 to 07:00 UT of day 63. Problems with a damaged DP.

Effelsberg (A. Polatidis at Tue Mar 7 16:16:40 2006) Minor failures :
Lost scans between 12:00-1600 UT and 2039-0023 UT due to snow

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)